# Multimodal Evaluation of Russian-language Architectures

### v2

@@ Abstract
Multimodal large language models (MLLMs) are currently at the center of research attention, showing rapid progress in scale and capabilities, yet their intelligence, limitations, and risks remain insufficiently understood. To address these issues, particularly in the context of the Russian language, where no multimodal benchmarks currently exist, we introduce MERA Multi, an open multimodal evaluation framework for Russian-spoken architectures. The benchmark is instruction-based and encompasses default text, image, audio, and video modalities, comprising 18 newly constructed evaluation tasks for both general-purpose models and modality-specific architectures (image-to-text, video-to-text, and audio-to-text). Our contributions include: (i) a universal taxonomy of multimodal abilities; (ii) 18 datasets created entirely from scratch with attention to Russian cultural and linguistic specificity, unified prompts, and metrics; (iii) baseline results for both closed-source and open-source models; (iv) a methodology for preventing benchmark leakage, including watermarking and licenses for private sets. While our current focus is on Russian, the proposed benchmark provides a replicable methodology for constructing multimodal benchmarks in typologically diverse languages, particularly within the Slavic language family.
@@
Artem Chervyakov*, Ulyana Isaeva*, Anton Emelyanov, Artem Safin, Maria Tikhonova, Alexander Kharitonov, Yulia Lyakh, Petr Surovtsev, Denis Shevelev, Vildan Saburov, Vasily Konovalov, Elisei Rykov, Ivan Sviridov, Amina Miftakhova, Ilseyar Alimova, Alexander Panchenko, Alexander Kapitanov, Alena Fenogenova* MERA Team Correspondence: mera@a-ai.ru
@@ 1 Introduction
Recent breakthroughs in generative AI, including models like GPT-5111https://openai.com/index/introducing-gpt-5, ImageBind girdhar2023imagebind, and LLaVa DBLP:conf/cvpr/LiuLLL24, have significantly advanced the state of the art across multiple modalities. This accelerated progress created a growing need for comprehensive multimodal benchmarks capable of rigorous evaluation of the full spectrum of versatile capabilities of such models. Although several benchmarks have been proposed for English and general-domain evaluation, such as MultiBench liang2021multibench, MMBench liu2024mmbench, and General-Bench fei2025path, they predominantly neglect the linguistic and cultural nuances of Slavic languages, particularly Russian. Beyond its Cyrillic script, Russian possesses a rich cultural context where concepts familiar to native speakers (e.g., folklore, Soviet media) are cognitively foreign to others, creating a challenge for automatic understanding.
Existing Russian-specific benchmarks, including TAPE taktasheva2022tape, Russian SuperGLUE fenogenova2022russian, and MERA fenogenova2024mera, focus exclusively on text-based tasks, leaving a critical gap in multimodal evaluation. To address this, we introduce MERA Multi, the first multimodal benchmark for MLLM evaluation in Russian. It comprises 18 tasks spanning (default) text [^2] Since textual input is an inherent component of all tasks in our benchmark, we treat it as the default modality. Consequently, we do not explicitly list it when describing or tabulating tasks, unless required for clarity., image, audio, and video modalities, built upon a unified taxonomy of multimodal abilities. Beyond Russian MERA Multi offers a blueprint for developing multimodal benchmarks across other Slavic and morphologically rich languages. These languages share structural complexity, typological proximity, and cultural specificity that make direct translation or adaptation of English-centric benchmarks inadequate. Thus, it serves not only as the first multimodal benchmark for Russian but also as a scalable methodology for culturally and linguistically aware evaluation in underrepresented settings.
@@
  • We create 18 novel datasets incorporating Russian cultural and linguistic specificities, unified prompts, and metrics.
@@
Additionally, we provide a standardized codebase for full reproducibility and a submission platform with automated scoring and a public leaderboard. [^3] https://mera.a-ai.ru/ru/multi All datasets and code are made available under the MERA Multi license, which allows the use of the benchmark sets for non-commercial purposes, provided that the data is not used for model training of any kind. These elements establish a foundation for transparent and culturally aware multimodal evaluation in Russian, while positioning MERA Multi as a reference point for developing similar frameworks across Slavic and other non-English languages, thereby promoting broader community development and cross-linguistic benchmarking.
@@ 2 Related Work
Evaluation of language models has historically relied on Natural Language Understanding benchmarks such as GLUE wang2018glue and SuperGLUE wang2019superglue, which set initial standards for English. As these benchmarks became saturated, newer instruction-oriented and reasoning-focused benchmarks emerged such as BIG-bench DBLP:conf/acl/SuzgunSSGTCCLCZ23 and HELM perlitz-etal-2024-efficient. Further efforts such as MMLU hendrycks2020measuring, AGIEval zhong2023agieval, and C-Eval huang2023c extended evaluation to academic and professional domains, though primarily in English and Chinese. For Russian, several text-based benchmarks were introduced. Among them are Russian SuperGLUE shavrina2020russiansuperglue; fenogenova2022russian, TAPE taktasheva2022tape, and RuCoLA mikhailov2022rucola. The instruction-based benchmark MERA fenogenova2024mera advanced this line of work but remained limited to text-only evaluation. Our work extends this direction toward multimodal, instruction-following assessment of MLLMs in Russian.
@@
The rapid progress of multimodal models has led to numerous benchmarks extending evaluation beyond text. Early large-scale efforts like MultiBench liang2021multibench covered 10 modalities and emphasized general-purpose representation learning. MMBench liu2024mmbench focused on fine-grained visual reasoning and bilingual (English–Chinese) evaluation, while SUPERB yang2021superb unified diverse audio tasks under a single framework. For video understanding, STAR wu2024star, InfiniBench ataallah2024infinibench, and Video-MME fu2025video advanced evaluation toward temporal and long-context reasoning. More comprehensive setups such as General-Bench fei2025path, OmniDialog razzhigaev2024omnidialog, MMMU yue2023mmmu, and SEED-Bench li2023seed assess multimodal and reasoning abilities at scale. All in all, despite these advances, most existing multimodal benchmarks are English-centric. To our knowledge, there is currently no existing multimodal benchmark for the Russian language. MERA Multi addresses the gap by providing a unified and culturally adapted evaluation of understanding across several modalities (see Appendix A for the comparison between MERA Multi and other benchmarks).
3 Overview of Multimodal Benchmark
@@
audio ruEnvAQA 596 0.0 MC CC BY-NC 4.0
RuSLUn 741 1.0 OE $C$-BY-4.0
*AQUARIA 738 0.0 MC MERA Multi
*ruTiE-Audio 1500 0 MC MERA Multi
image ruCLEVR 1148 73.8 OE $C$-BY-4.0
ruCommonVQA 3015 1.0 OE $C$-BY-4.0
ruNaturalScienceVQA 363 70.0 MC $C$-BY-SA-4.0
WEIRD 814 82.22 MC $C$-BY-4.0
*LabTabVQA 339 93.9 MC MERA Multi
*RealVQA 773 1.0 OE MERA Multi
*ruHHH-Image 595 0 MC MERA Multi
*ruMathVQA 502 81.0 OE MERA Multi
*ruTiE-Image 1500 0 MC MERA Multi
*SchoolScienceVQA 4227 52.0 MC MERA Multi
*UniScienceVQA 7432 1.0 OE MERA Multi
video CommonVideoQA 1200 95.75 MC $C$-BY-4.0
*RealVideoQA 671 95.83 MC MERA Multi
*ruHHH-Video 911 0 / 0 MC MERA Multi
Table 1: Overview of datasets in MERA Multi. Those marked with an asterisk are private datasets collected from scratch, while the others are public datasets compiled from open-source datasets. Size column shows the number of samples in the dataset, and Answer column is the task format (MC and OE stand for multiple-choice and open-ended, respectively). HB is the human baseline value (basic / expert or basic only, see G.2 for details). MERA Multi license refers to the benchmark license anonimized for the review period.
@@ 3.1 Benchmark General Structure
To balance reproducibility and novelty, the benchmark integrates both publicly available datasets (7 tasks), curated from open data sources, and private datasets (11 tasks), collected from scratch specifically for this study. The latter are designed to incorporate Russian cultural nuances, target underexplored skill categories, and mitigate potential data contamination issues (see details in section 3.4). A complete list of datasets is provided in Table 1. Task examples and dataset creation details are in B.
@@ 3.2 Skill Taxonomy
Contemporary studies on multimodal benchmarks often rely on custom skill sets during the dataset design process liu2024mmbench; fu2024mmecomprehensiveevaluationbenchmark. We performed a comprehensive analysis of such systems and synthesized them into a consolidated MLLM skill taxonomy, which underpins the foundation of MERA Multi. Such a system functions as a comprehensive map for skill coverage, which, when coupled with an alignment of existing datasets to corresponding skills, can spotlight deficiencies in benchmark task diversity.
The taxonomy is aligned with three broad categories of human-like cognition: perception, knowledge, and reasoning, a division also adopted by recent multimodal benchmarks. The knowledge taxonomy is shown in Table 2, and the perception and reasoning parts are provided in the Appendix in Table 7 and Table 8 respectively, due to space constraints. To comply with the emerging MLLM capabilities, this taxonomy is designed to be extendable and this paper provides the initial version of the unified taxonomy and encourages its adoption and extension in future research.
@@
Further details on skill taxonomy are in Appendix C.
@@ 3.3.1 Evaluation Framework Implementation
We build the benchmark code base on the lm-eval [^4] https://github.com/EleutherAI/lm-evaluation-harness framework (eval-harness; eval-harness-paper), extending it to support multimodal inputs while preserving its core structure for texts. The codebase introduces comprehensive vision, audio, and video evaluation through, and chat-template formatting tailored for instruction-tuned and API-based models. Multimodal data is integrated either by passing it separately to the model’s processor or by embedding it directly within a chat template alongside the text. All evaluations are strictly generative: models produce free-form answers until a stop condition is met. All outputs are used directly except RuSLUn which requires minimal parsing for structured output.
@@ 3.3.2 Prompt Structure
Following the approach of voronov2024mind, we avoid hard-coding task-specific prompts by employing a block-prompting scheme that constructs universal prompts aligned with each task’s structure. For every dataset, we instantiate a predefined set of 10 prompts drawn from fixed block layouts that we initially designed and realized under different surface modalities (e.g., formal request, command, technical statement). This design preserves a uniform formulation across benchmark tasks while allowing for controlled variation, such as the presence or absence of a reasoning request, to mitigate benchmark saturation.
Following the methodology of fenogenova2024mera, we uniformly assign prompts for each task across dataset samples to ensure that the aggregated performance metric reflects an average over prompt variants rather than the idiosyncrasies of a single prompt. A post-hoc statistical analysis (see Appendix F for the details) confirmed that our prompt set is not invariant, as some prompts produce statistically significant shifts in model scores. This finding directly motivates and validates our multi-prompt design as a necessary guard against evaluation bias.
@@ 3.3.3 Judge Scoring
This judge is trained on a diverse, human-annotated dataset of model outputs collected from various Russian-language instruction-based benchmarks. The final model, based on the RuModernBERT [^5] https://hf.co/deepvk/RuModernBERT-base deepvk2025rumodernbert encoder, achieves F1 score of 0.96 on a held-out test set and shows 99.6% agreement with EM on identical answers, confirming its reliability. We deploy this judge as our primary metric, providing a robust, scalable score for correctness across the benchmark. Comprehensive details on data, training, and model comparisons are in the Appendix E.
@@ 3.4 Data Protection
As pre-training datasets grow, benchmark leakage is becoming more common. This issue is exacerbated by opaque training processes and the undisclosed use of supervised data in modern MLLMs, which undermines the validity of benchmarks and fosters unfair comparisons, ultimately slowing progress. To protect the multimodal data in our private benchmark suite from unauthorized use, we employ a three-part strategy: (i) data watermarking (Section 3.4.1), (ii) leakage detection (Section 3.4.2), and (iii) licensing (Section 3.4.3). Together, these measures provide robust technical and legal safeguards.

### v3

@@ Abstract
Multimodal large language models (MLLMs) are currently at the center of research attention, showing rapid progress in scale and capabilities, yet their intelligence, limitations, and risks remain insufficiently understood. To address these issues, particularly in the context of the Russian language, where no multimodal benchmarks currently exist, we introduce MERA Multi, an open multimodal evaluation framework for Russian-spoken architectures. The benchmark is instruction-based and encompasses default text, image, audio, and video modalities, comprising 18 newly constructed evaluation tasks for both general-purpose models and modality-specific architectures (image-to-text, video-to-text, and audio-to-text). Our contributions include: (i) a universal taxonomy of multimodal abilities; (ii) 18 datasets created entirely from scratch with attention to Russian cultural and linguistic specificity, unified prompts, and metrics; (iii) baseline results for both closed-source and open-source models; (iv) a methodology for preventing benchmark leakage, including watermarking for private sets. While our current focus is on Russian, the proposed benchmark provides a replicable methodology for constructing multimodal benchmarks in typologically diverse languages, particularly within the Slavic language family.
@@
Artem Chervyakov*1, Ulyana Isaeva*1, Anton Emelyanov1, Artem Safin1, Maria Tikhonova1,4, Alexander Kharitonov1, Yulia Lyakh1, Petr Surovtsev1, Denis Shevelev1, Vildan Saburov1,8, Vasily Konovalov3,5,6, Elisei Rykov5,7, Ivan Sviridov2, Amina Miftakhova2, Ilseyar Alimova5, Alexander Panchenko5,3, Alexander Kapitanov1, Alena Fenogenova*1 1SberAI, 2Sber AI Lab, 3AIRI, 4HSE University, 5Skoltech, 6MIRAI, 7T-Tech, 8Moscow Center for Advanced Studies Correspondence: mera@a-ai.ru
@@ 1 Introduction
Recent breakthroughs in generative AI, including models like GPT-5111https://openai.com/index/introducing-gpt-5, ImageBind Girdhar et al. (2023), and LLaVa Liu et al. (2024a), have significantly advanced the state of the art across multiple modalities. This accelerated progress created a growing need for comprehensive multimodal benchmarks capable of rigorous evaluation of the full spectrum of versatile capabilities of such models. Although several benchmarks have been proposed for English and general-domain evaluation, such as MultiBench Liang et al. (2021), MMBench Liu et al. (2024e), and General-Bench Fei et al. (2025), they predominantly neglect the linguistic and cultural nuances of Slavic languages, particularly Russian. Beyond its Cyrillic script, Russian possesses a rich cultural context where concepts familiar to native speakers (e.g., folklore, Soviet media) are foreign to others, creating a challenge for automatic understanding.
Existing Russian-specific benchmarks, including TAPE Taktasheva et al. (2022), Russian SuperGLUE Fenogenova et al. (2022), and MERA Fenogenova et al. (2024), focus exclusively on text-based tasks, leaving a critical gap in multimodal evaluation. To address this, we introduce MERA Multi [^2] https://mera.a-ai.ru/en/multi, the first multimodal benchmark for MLLM evaluation in Russian. It comprises 18 tasks spanning (default) text [^3] Since textual input is an inherent component of all tasks in our benchmark, we treat it as the default modality. Consequently, we do not explicitly list it when describing or tabulating tasks, unless required for clarity., image, audio, and video modalities, built upon a unified taxonomy of multimodal abilities. Beyond Russian MERA Multi offers a blueprint for developing multimodal benchmarks across other Slavic and morphologically rich languages. These languages share structural complexity, typological proximity, and cultural specificity that make direct translation or adaptation of English-centric benchmarks inadequate. Thus, it serves not only as the first multimodal benchmark for Russian, but also as a scalable methodology for culturally and linguistically aware evaluation in underrepresented settings.
@@
  • We create 18 novel datasets  [^4] https://hf.co/collections/MERA-evaluation/mera-multimodality incorporating Russian cultural and linguistic specificities, unified prompts, and metrics.
@@
Additionally, we provide a standardized codebase for full reproducibility and a submission platform with automated scoring and a public leaderboard [^5] https://mera.a-ai.ru/en/multi/leaderboard. All datasets and code are made available under the MERA Multi license, which allows the use of the benchmark sets for non-commercial purposes, provided that the data is not used for model training of any kind. These elements establish a foundation for transparent and culturally aware multimodal evaluation in Russian, while positioning MERA Multi as a reference point for developing similar frameworks across Slavic and other non-English languages, thereby promoting broader community development and cross-linguistic benchmarking.
Benchmark Modalities # Tasks/Skills Primary Focus Language Cultural Focus
MultiBench Liang et al. (2021) 10 (Text, Image, Audio, Video, Table, Set, …) 20 General-purpose representation EN General
MMBench Liu et al. (2024e) Image, Text 20 Fine-grained visual reasoning EN, ZH General
SUPERB Yang et al. (2021) Speech, Audio 10 Speech processing EN General
STAR Wu et al. (2021) Video, Text 4 VideoQA, temporal reasoning EN General
InfiniBench Ataallah et al. (2024) Video, Text 8 Long-context video understanding EN General
Video-MME Fu et al. (2025) Video, Text 12 Fine-grained video analysis EN General
General-Bench Fei et al. (2025) Text, Image, Video, Audio, Tabule 700+ Large-scale ability coverage EN General
OmniDialog Razzhigaev et al. (2024) Text, Image, Audio 8 Multimodal dialogue EN General
MMMU Yue et al. (2024) Image, Text 30 Expert-level reasoning EN General
SEED-Bench Li et al. (2023) Image, Video, Text 12 Generative comprehension EN General
MERA Multi (ours) Text, Image, Audio, Video 18 Comprehensive understanding RU General + Russian
@@ 2 Related Work
Evaluation of language models has historically relied on Natural Language Understanding benchmarks such as GLUE Wang et al. (2019b) and SuperGLUE Wang et al. (2019a), which set initial standards for English. As these benchmarks became saturated, newer instruction-oriented and reasoning-focused benchmarks emerged such as BIG-bench Suzgun et al. (2023) and HELM Perlitz et al. (2024). Further efforts such as MMLU Hendrycks et al. (2021), AGIEval Zhong et al. (2024), and C-Eval Huang et al. (2023) extended evaluation to academic and professional domains, though primarily in English and Chinese. For Russian, several text-based benchmarks were introduced. Among them are Russian SuperGLUE Shavrina et al. (2020); Fenogenova et al. (2022), TAPE Taktasheva et al. (2022), and RuCoLA Mikhailov et al. (2022). The instruction-based benchmark MERA Fenogenova et al. (2024) advanced this line of work but remained limited to text-only evaluation. Our work extends this direction toward multimodal, instruction-following assessment of MLLMs in Russian.
@@
The rapid progress of multimodal models has led to numerous benchmarks extending evaluation beyond text. Early large-scale efforts as MultiBench Liang et al. (2021) covered 10 modalities and emphasized general-purpose representation learning. MMBench Liu et al. (2024e) focused on fine-grained visual reasoning and bilingual (English–Chinese) evaluation, while SUPERB Yang et al. (2021) unified diverse audio tasks under a single framework. For video understanding, STAR Wu et al. (2021), InfiniBench Ataallah et al. (2024), and Video-MME Fu et al. (2025) advanced evaluation toward temporal and long-context reasoning. More comprehensive setups such as General-Bench Fei et al. (2025), OmniDialog Razzhigaev et al. (2024), MMMU Yue et al. (2024), and SEED-Bench Li et al. (2023) assess multimodal and reasoning abilities at scale. All in all, despite these advances, most existing multimodal benchmarks are English-centric. To our knowledge, there is currently no existing multimodal benchmark for the Russian language. MERA Multi addresses the gap by providing a unified and culturally adapted evaluation of understanding across several modalities (see Table 1 for the detailed comparison between MERA Multi and other benchmarks).
3 Overview of MERA Multi Benchmark
@@
audio ruEnvAQA 596 0.95 MC CC BY-NC 4.0
RuSLUn 741 0.90 OE $C$-BY-4.0
*AQUARIA 738 0.98 MC MERA Multi
*ruTiE-Audio 1500 0.75 MC MERA Multi
image ruCLEVR 1148 0.96 OE $C$-BY-4.0
ruCommonVQA 3015 0.84 OE $C$-BY-4.0
ruNaturalScienceVQA 363 0.99 MC $C$-BY-SA-4.0
WEIRD 814 0.85 MC $C$-BY-4.0
*LabTabVQA 339 0.91 MC MERA Multi
*RealVQA 773 0.63 OE MERA Multi
*ruHHH-Image 595 0.89 MC MERA Multi
*ruMathVQA 502 0.95 OE MERA Multi
*ruTiE-Image 1500 0.77 MC MERA Multi
*SchoolScienceVQA 4227 0.82 MC MERA Multi
*UniScienceVQA 7432 0.13 OE MERA Multi
video CommonVideoQA 1200 0.96 MC $C$-BY-4.0
*RealVideoQA 671 0.96 MC MERA Multi
*ruHHH-Video 911 0.84 MC MERA Multi
Table 2: Overview of datasets in MERA Multi. Those marked with an asterisk are private datasets collected from scratch, while the others are public datasets compiled from open-source datasets. Size column shows the number of samples in the dataset. HB is the human baseline value (basic / expert or basic only, see F.2 for details). Answer column is the task format (MC and OE stand for multiple-choice and open-ended, respectively). MERA Multi license refers to the benchmark license anonymized for the review period.
Figure 1 presents an overview of MERA Multi and its methodological parts.
@@ 3.1 Benchmark General Structure
To balance reproducibility and novelty, the benchmark integrates both publicly available datasets (7 tasks), curated from open data sources, and private datasets (11 tasks), collected from scratch specifically for this study. The latter are designed to incorporate Russian cultural nuances, target underexplored skill categories, and mitigate potential data contamination issues (see details in section 3.4). A complete list of datasets is provided in Table 2. Task examples and dataset creation details are in A.
@@ 3.2 Skill Taxonomy
Contemporary studies on multimodal benchmarks often rely on custom skill sets during the dataset design process Liu et al. (2024e); Fu et al. (2024). We performed a comprehensive analysis of such systems and synthesized them into a consolidated MLLM skill taxonomy, which underpins the foundation of MERA Multi. Such a system functions as a comprehensive map for skill coverage, which, when coupled with an alignment of existing datasets to corresponding skills, can spotlight deficiencies in benchmark task diversity.
The taxonomy is aligned with three broad categories of human-like cognition: knowledge, perception, and reasoning, a division also adopted by recent multimodal benchmarks. The knowledge taxonomy is shown in Table 3, and the perception and reasoning parts are provided in the Appendix in Table 7 and Table 8 respectively, due to space constraints. To comply with the emerging MLLM capabilities, this taxonomy is designed to be extendable and this paper provides the initial version of the unified taxonomy and encourages its adoption and extension in future research.
@@
Further details on skill taxonomy are in Appendix B.
@@ 3.3.1 Evaluation Framework Implementation
We build the benchmark code base on the lm-eval [^6] https://github.com/EleutherAI/lm-evaluation-harness framework (Gao et al., 2024; Biderman et al., 2024), extending it to support multimodal inputs while preserving its core structure for texts. The codebase introduces comprehensive vision, audio, and video evaluation through, and chat-template formatting tailored for instruction-tuned and API-based models. Multimodal data is integrated either by passing it separately to the model’s processor or by embedding it directly within a chat template alongside the text. All evaluations are strictly generative: models produce free-form answers until a stop condition is met. All outputs are used directly except RuSLUn which requires minimal parsing for structured output.
@@ 3.3.2 Prompt Structure
Following the approach of Voronov et al. (2024), we avoid hard-coding task-specific prompts by employing a block-prompting scheme that constructs universal prompts aligned with each task’s structure. For every dataset, we instantiate a predefined set of 10 prompts drawn from fixed block layouts that we initially designed and realized under different surface modalities (e.g., formal request, command, technical statement). This design preserves a uniform formulation across benchmark tasks while allowing for controlled variation, such as the presence or absence of a reasoning request, to mitigate benchmark saturation.
Following the methodology of Fenogenova et al. (2024), we uniformly assign prompts for each task across dataset samples to ensure that the aggregated performance metric reflects an average over prompt variants rather than the idiosyncrasies of a single prompt. A post-hoc statistical analysis (see Appendix E for the details) confirmed that our prompt set is not invariant, as some prompts produce statistically significant shifts in model scores. This finding directly motivates and validates our multi-prompt design as a necessary guard against evaluation bias.
@@ 3.3.3 Judge Scoring
This judge is trained on a diverse, human-annotated dataset of model outputs collected from various Russian-language instruction-based benchmarks. The final model, based on the RuModernBERT [^7] https://hf.co/deepvk/RuModernBERT-base Spirin et al. (2025) encoder, achieves F1 score of 0.96 on a held-out test set and shows 99.6% agreement with EM on identical answers, confirming its reliability. We deploy this judge as our primary metric, providing a robust, scalable score for correctness across the benchmark. Comprehensive details on data, training, and model comparisons are in the Appendix D.
@@ 3.4 Data Protection
As pre-training datasets grow, benchmark leakage is becoming more common. This issue is exacerbated by opaque training processes and the undisclosed use of supervised data in modern MLLMs, which undermines the validity of benchmarks and fosters unfair comparisons, ultimately slowing progress. To protect the multimodal data in our private benchmark suite from unauthorized use, we employ: (i) data watermarking (Section 3.4.1), (ii) leakage detection (Section 3.4.2). We also introduce the license, which explicitly permits the use of the benchmark data for research and non-commercial testing purposes, but strictly prohibits its incorporation into any model training process. Together, these measures provide robust technical and legal safeguards.